OpenStudy (anonymous):

Determine the factored form of the following polynomial: 4x^2+33x+35

OpenStudy (anonymous):

(4x + 5) (x + 7)

OpenStudy (anonymous):

(4x + 5)(x +7)

OpenStudy (owlfred):

Come on guys, instead of just the answer, let's explain why.

OpenStudy (anonymous):

ok ..

OpenStudy (anonymous):

watch me Lyssah

OpenStudy (anonymous):

Kay .

OpenStudy (amistre64):

and, one way to curb unethical behaviour as well as to provide just enough information for them to reach the answer without actually giving them the answer :)

OpenStudy (amistre64):

..as well is to .....

OpenStudy (anonymous):

\[4x^2 + 33x + 35=0\] Now if I split it into 7x and 5x \[4x^2 + 28x + 5x + 35=0\] \[4x(x + 7) +5(x+7)=0\] Now all you have to do is take x+7 common

OpenStudy (anonymous):

thanks owlfred

OpenStudy (anonymous):

lyssah to thankyou well did you get that?

OpenStudy (anonymous):

oh sorry i made a typo ther i have to split 33x into 28x and 5x